Suppose that y varies inversely with x. When x = 10​, then y = 9. Find y when x = 5.
​(a) Find the constant of proportionality k.
​(b) Write the function relating the variables.
​(c) Find the indicated quantity.

Answer by Boreal(15235)

y=k/x
9=k/10
k=90
y=90/5
=18.